OpenStudy (anonymous):

Is anybody out there good with circuits? Kirchhoff's rules are killing me. I'm probably over thinking it but I am in dire need of some external resources that will help the concept sink in.

OpenStudy (anonymous):

what do u need to understand?

OpenStudy (anonymous):

its basically a law of nodes, u need to know how many and where the nodes are of each circuit, and the law is the sum of the currents entering each node is equal to the sum of the currents going out of the node,

OpenStudy (anonymous):

 So this is how I labeled my circuit. The arrows indicate the direction of current moving through each resistor. From what I understand, there is no "wrong way" to label the direction as long as I understand that if I get a negative value for potential (V=IR), it implies that I was going the "wrong way" or against the current instead of with it. I kind of understand the node/junction rule. But I'm not sure if I am applying it correctly. Am I justified in thinking that: \[I4=I1+I2\] and \[I3=I1+I4\]

OpenStudy (radar):

Do not understand the need for both I3 and I2 as they are both the same

OpenStudy (anonymous):

@radar how do you know are they the same?

OpenStudy (anonymous):

At a second glance I thought maybe \[I3=I4-I1 \] Which would equal I2.... Is that right?

OpenStudy (radar):

The current thru R3 is the same as the current thru R2 as R3 and R2 are connected in series with each other. Although that series arrangement of R2 and R3 are parallel with R4, this will not alter the fact that the current thru R2 must also go thru R3.

OpenStudy (radar):